P12 A soccer ball is kicked such that it barely clears a 3 m tall fence located 40 m away. If it took the ball 3 s to go from the point where it was struck to the fence, a) b) c) What was the initial velocity of the ball? What is the position of the ball when it is at the maximum height? What is the velocity of the ball when it hits the ground, after going past the fence? Kicker 3 m 40 m
Answer #1

a)

Initial Horizontal and vertical components of velocities are

V_{ox}=V_{o}Cos heta

V_{oy}=V_{o}Sin heta

Horizontal distance traveled

X=V_{ox}t

40 -Vor13.33m/s

From

Y=Y_{o}+V_{oy}t-rac{1}{2}gt^{2}

30+Voy(3)- (9.81)(3)2

Voy-15.715m/s15.72m/s

overrightarrow{V_{o}}=(13.33widehat{i}+15.72widehat{j})m/s

b)

From

V_{fy}=V_{oy}-gt

at maximum height Vfy=0

0=15.715-9.81*t

t=1.602 s

From

V_{fy}^{2}=V_{oy}^{2}-2gh

0=15.7152-2*9.81*H

H=12.587 m=12.6 m

Horizontal position

XH=1.602*13.333=21.36 m=21.4 m

overrightarrow{R}=(21.4widehat{i}+12.6widehat{j})m

c)

when the object hits ground Y=0 ,from

Y=Yo+Voyt-(1/2)gt2

0=0+15.715*t-(1/2)*9.81*t2

t=3.204 s

Vfx=Vox=13.333 m/s=13.33 m/s

Vfy=Voy-gt =15.715-9.81*3.204

Vfy=-15.715 m/s=-15.72 m/s

у,-(13.33ì-15.72)m/s 2 29